OpenStudy (nicole143):

@mathmale Hello, could you help me with finding all the zeros of the function? y= (x+3)(x+2)(x-2) (Is it -3, -2, 2?)

OpenStudy (mathmale):

Hi, N, I'd be happy to help. At the moment I'm working with another student. I'll get back to you asap, but encourage you to "hear" what others on Open Study have to contribute also.

OpenStudy (nicole143):

@Saphira_Shapiro Huh? @mathmale Okay, thank you!

OpenStudy (unklerhaukus):

@Nicole143 you done it right, for the left hand side to be zero, at least one of the factors on the right hand side must be zero so you did (x+3)=0 or (x+2)=0 or (x-2)=0 which you simplified to get x= -3, -2, or 2 √√√

OpenStudy (nicole143):

Awesome, could you help me with another? And would you like me to open it in another question?

OpenStudy (unklerhaukus):

sure,

OpenStudy (nicole143):

@UnkleRhaukus Okay, so I have to find all the zeros of the equation 27x^2 - 324 = -x^4

OpenStudy (unklerhaukus):

first get all the terms on the right hand side, then i would substitute x^2=z, x^4=z^2 to get a quadratic

OpenStudy (nicole143):

Okay, I now have 27z + z^2 - 324 = 0

OpenStudy (unklerhaukus):

can you factor that , with the quadratic equation?

OpenStudy (nicole143):

With the equation x = -b+- the square root of b^2 - 4ac / 2a?

OpenStudy (unklerhaukus):

yeah that's the quadratic formula, can you get the factors now?

OpenStudy (nicole143):

I don't know.. I don't think it's quite clicking.

OpenStudy (unklerhaukus):

actually you dont need to explicitly find the factors, the quadratic formula will give you the zeros straight away

OpenStudy (nicole143):

So just put in 27z for a, z^2 for b, and -324 for c?

OpenStudy (unklerhaukus):

z^2+27z - 324 = 0 a=1 b=27 c=-324

OpenStudy (nicole143):

Oh, okay. One moment.

OpenStudy (nicole143):

I now have

OpenStudy (nicole143):

I didn't hit post:/ I have x = -27 +- the square root of 2592 / 2

OpenStudy (nicole143):

Or are you supposed to leave for as a positive to be able to minus?

OpenStudy (unklerhaukus):

z=[-b±√(b^2-4ac)]/2a =[-27±√(27^2-4(1)(-324))]/2 =[-27±√(2025)]/2

OpenStudy (mathmale):

Are you still working on the original problem ("Hello, could you help me with finding all the zeros of the function? y= (x+3)(x+2)(x-2) )? Need me, or are you two getting where you want to go?

OpenStudy (unklerhaukus):

we done that one already @mathmale

OpenStudy (nicole143):

I would love for you to help as well - I'm a bit off on my answer at least for what he has. And yes we are on a different one now.

OpenStudy (mathmale):

Great! Nicole, I need to get off the computer very soon. But you have my address in case you want to share future problems with me.

OpenStudy (nicole143):

I unfortunately need to get this done tonight but hopefully someone will stick around and help(: @mathmale

OpenStudy (mathmale):

I see that U-R has explained some things quite clearly and competently. What specific part of your discussion is not yet clear for you?

OpenStudy (nicole143):

Will you be around long enough to explain a different question to me? @mathmale For this one I just didn't get the sam answer as him..

OpenStudy (nicole143):

*same

OpenStudy (unklerhaukus):

i kinda have to go now, so you can help with this second one @mathmale dont forget when you get the z's you still need to get the x's by taking ±sqrts

OpenStudy (mathmale):

Nicole, I am preparing to leave tomorrow morning on a 1,500 mile trip to Canada. However, if you'll post the next question I'll get back to you after a few minutes away from my computer. Thank you very much U-R!

OpenStudy (nicole143):

Okay, thank you for your help! @UnkleRhaukus

OpenStudy (nicole143):

@mathmale Okay(:

OpenStudy (mathmale):

Nicole, either post your question here, or email it to me. Your choice. Or, send it me as a message.

OpenStudy (nicole143):

I'll email you.

OpenStudy (nicole143):

@mathmale I'll send it in a few minuets, thank you for the help!